Vilas Company is considering a capital investment of $216,000 in additional productive facilities. The new machinery is expected to have a useful life of 5 years with no salvage value. Depreciation is by the straight-line method. During the life of the investment, annual net income and net annual cash flows are expected to be $18,468 and $45,000, respectively. Vilas has a 12% cost of capital rate, which is the required rate of return on the investment.

Click here to view PV table.

(a)

Compute the cash payback period. (Round answer to 1 decimal place, e.g. 10.5.)

Cash payback period

years


Compute the annual rate of return on the proposed capital expenditure. (Round answer to 2 decimal places, e.g. 10.52%.)

Annual rate of return

%


(b)

Using the discounted cash flow technique, compute the net present value. (If the net present value is negative, use either a negative sign preceding the number e.g. -45 or parentheses e.g. (45). Round answer for present value to 0 decimal places, e.g. 125. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Net present value
Answer #1

a) cash payback period = Initial investment/Annual Cash flow = 216000/45000 = 4.8 years

Annual rate of return = 18468/216000 = 8.55%

b) Net present value = (45000*3.60478)-216000 = -53785